OpenStudy (anonymous):

How many 2 over 3 foot pieces can Wendy cut from the 5 and 1 over 3 feet of ribbon? Show your work. (5 points)

OpenStudy (anonymous):

will fan and medal

OpenStudy (anonymous):

@marylou004 ,@dumbcow

OpenStudy (anonymous):

@dumbcow

OpenStudy (anonymous):

@dolphincove

OpenStudy (anonymous):

HELP ME PLZ

OpenStudy (anonymous):

And you simplify the question? "1 over 3 feet" doesn't make much sense...

OpenStudy (anonymous):

Can*

OpenStudy (anonymous):

hu

OpenStudy (anonymous):

'hu' isn't a word... can you speak english please?

OpenStudy (anonymous):

yeah i can speak english and do u have a problem with me saying hu

OpenStudy (anonymous):

@Shroomsy

OpenStudy (dumbcow):

just divide \[\frac{ 5 \frac{1}{3}}{\frac{2}{3}} = \frac{16}{3}*\frac{3}{2}\]

OpenStudy (anonymous):

divide what

OpenStudy (dumbcow):

total length of ribbon divided by desired piece length

OpenStudy (anonymous):

soo 5 and 1 3rd divided by 2/3

OpenStudy (anonymous):

@dumbcow

OpenStudy (anonymous):

@uri

OpenStudy (anonymous):

can u help me

OpenStudy (uri):